OpenStudy (anonymous):

Whenever Art Dealer sells a sculpture, he earns a 20 percent commission on the first $12,000 of the sale price plus 15 percent of the sale price in excess of $12,000. If Art earned a $3,900 commission on the sale of a certain sculpture, what was the sale price?

OpenStudy (manish00333):

20% of 12,000 = 2,400 15% = 0.15 3,900 = 2,400 + 0.15x 1,500 = 0.15x x = 10,000 = sale price in excess of $12,000 so, 10, 000 + 12, 000 = $22, 000